In a rectangle, the perimter is 94 inches. The length of the rectangle is 5 inches less than three times the width. What is the length and width?
Found 2 solutions by mananth, solver91311:
Answer by mananth(16949)

let width =x
length = 3x-5
perimeter =94 inches
..
Perimeter = 2L + 2W
2x+2(3x-5)=94
2x+6x-10=94
8x=104
x=13 inches the width
length = 3*13-5
=34 inches
..
2*13+2*34= 94

Answer by solver91311(24713)



Let represent the width. Then the length must be: . The perimeter is two times the length plus two times the width.



Solve for

John